OpenStudy (anonymous):

Golf clubs R Us had a set of 10 gold clubs that were marked on sale for $510. This was a discount off the original selling priceIf the golf clubs cost Gold Clubs R Us $490, what was their profit?

OpenStudy (anonymous):

If you will, help me figure out how to do it on my own :)

OpenStudy (btaylor):

If you buy something for $100 and plan to sell it to someone else, you want to make some money off of it. So, you sell it for $110, and pocket the $10 profit. Profit = Price you sold it for - Price you paid for it. In your situation, they paid $490, and sold it for $510. So, the profit = 510-490 = $20. An interesting side note: Businesses are very interested in profit, since that is how they manage to make money and stay in business. Except on Black Friday, when some stores sell goods for a loss, just so they can get rid of them.

OpenStudy (anonymous):

Hey, thanks so much BTaylor! That is a very interesting fact, btw!

OpenStudy (anonymous):

Oh there's more, if you're willing to help a bit longer :)

OpenStudy (btaylor):

you're welcome. I'm available and willing. :) Another trivia fact: In some parts of Oklahoma, it was illegal from 1935-2012 to sell goods for less than you paid for them. So this year was the first time people in some areas could shop on Black Friday.

OpenStudy (anonymous):

Golf clubs R Us had a set of 10 gold clubs that were marked on sale for $510. This was a discount off 10% of the original selling price Wgat was the store's percent of profit based on const? (490) Follow the problem-solving process and round your answer to the nearest 100th if necessary.

OpenStudy (anonymous):

Wow that's really cool! I've never heard that before. Are you a history freak too? That's what I call myself, anyway.

OpenStudy (btaylor):

So, the percent profit is the profit divided by the original price (here, 490). So, we have (510-490)/(490) = 20/490 = 4.08%

OpenStudy (anonymous):

Oh my gosh, someone finally explained this to me in terms I could understand! (excuse all the typos above, keyboard had a freakout)

OpenStudy (btaylor):

And no, I just heard that on the radio.